04/30/04 9:00 12m If you are getting ssh errors trying to log into modelo, then I need to remind you about the fact that we have a new computer called modelo. It is a Linux box replacing the old Sun. The error will look something like this. This is perfectly ok and you will need to remove any lines in your  ~/.ssh/known_hosts that contain the word modelo. News
40/27/04 9:00 12m The 1mm receiver is being warmed up to replace the refrigerator. Ongoing
04/27/04 9:00 12m The network to the 12 Meter seems to be working better, but a test of the throughput packet size indicates that any packet larger then 1500 bytes fails to get through. ??? Ongoing
04/26/04 17:30 12m We are experiencing network troubles between the 12 Meter and the top of Kitt Peak. Ongoing
04/26/04 16:45 12m The 1mm receiver has warmed up and is out of service at this time. Ongoing
04/21/04 3:30 12m The north South translation stage has been disabled and locked into the correct position. This feature was never used anyway and only caused problems. Fixed
04/21/04 3:30 12m The #2 UPS lost another battery (3B) and was replaced with a spare(1S). This explains the computer outages of this AM. Fixed
04/20/04 11:30 12m Halleluiah, e-mail to Kitt Peak finally works. Send any and all emails for the operator to: oper12m{at}gmail{dot}com Fixed
04/20/04 4:30 12m The 23mm receiver is once again cold. We will test it later today and evaluate the state of the mixers.  
04/19/04 9:00 12m The observers computer, modelo, has been reinstalled. This is the first in a multi-step program to replace all the computers on Kitt Peak with Linux based Intel systems. For a more complete description and how to survive during this transition period, see the 12 Meter Computer Conversion Document here. News
04/16/04 2:30 ARO We now feature DVD and CDROM formats for data archiving. This will be our primary method of data delivery. News
04/16/04 2:00 ARO The cookbooks, manuals and example books for our analysis packages are now on line. Click here to view. News
04/15/04 18:00 SMT Once again the network to Mt. Graham stopped working. A call to CCIT and they 'fixed' it again. Network was out for 2 hours. Ongoing
04/15/04   12m The new modelo has been installed and massive restructuring of the file systems is being undertaken. This means there may be things we overlooked. For the time being, continue to use locura for you observing needs. I will announce here when modelo is ready for prime-time. Ongoing
04/13/04 14:38 SMT The network problems to Mt. Graham have reportedly been corrected by CCIT. Fixed
04/12/04 15:00 12m The RED 2 MHz filter banks have been reinstalled. Fixed
04/12/04 14:00 12m 6 Batteries were replaced in the two Nova UPS's and the entire system load tested. All systems appear to function correctly now. Fixed
04/09/04   SMT The DesertStar receiver was used tonight utilizing all three mixers on the sky. A complete map of OriIrc2 was obtained in under 10 hours. Complete focusing, pointing and beam maps were also obtained. News
04/09/04 14:30 12m The MAC Appears to work properly now. Fixed
04/09/04 13:45 SMT The ongoing saga of the lousy network to Mt. Graham continues. Apparently, installing the new gateway, at the base camp, only solved the poor network issues temporarily. CCIT is still looking into the problem. Ongoing
04/09/04 13:20 ARO The new observers computer, 'modelo', has been purchased, configured to the first order and will be installed on Kitt Peak on Thursday the 15th. Here are some of its specs:
  1. Dual 2.66 GHz Xeon Processors.
  2. 1 GByte of memory.
  3. 360 GBytes of external disk space.
  4. Twin 19" flat panel LCD video displays.
  5. DVD-RW, CD-R, CDROM drive.
  6. Best of all, it's running Linux!!!

Observer software:

  1. Aips -- After running the DDT and Y2K tests, produces AipsMarks in the range of 118 - 125. I'm told that's extremely good. There are 5 aips disks worth 36 Gbytes each.
  2. Class -- being installed as I write this. There are (5) 36 Gbyte disks for reduced class data.
  3. Unipops -- In the process of being converted to Linux and will be installed when it's ready. In the meanwhile observers will need to login to locura and run unipops there, with their DISPLAY set back to 'modelo:0.0'.
News
04/09/04 11:05 12m The issue with the VME controller on the MAC was caused by a adjacent VME card not fully seated. Now lets hope things will work once the filterbank observing stops long enough to test the MAC. Fixed
04/09/09 11:00 12m It's entirely possible that I've managed to fix the outgoing mail. Now to get the incoming mail to the operators working. Ongoing
04/09/04 8:10 SMT The 19 Channel bolometer is cold and ready now.  
04/09/04 8:00 SMT The 345 sis is cooling down now.  
04/08/04 21:35 12m The 23mm receiver is now cold and working. Fixed
04/08/04 16:40 SMT The gateway at the MGIO base camp was replaced in an attempt to solve the network problems to Mt. Graham. Ongoing
04/08/04 14:00 12m There was a power failure on site that the UPS's did not cover, so we lost the computers. It appears that the MAC is having issues with one of the FFT computers, bigmac1. It seems to have forgotten it's a VME computer and is refusing to load the VME drivers. Fixed
04/10/09 10:00 SMT The Sun avoidance zone has been increased to 35 degrees in response to overheating of the feedlegs. News
04/02/04 14:25 12m As promised, in the remote observers manual, telnet has been dropped as a supported means of communication to Kitt Peak. "ssh" is the only way to log into out computers on Kitt Peak. News
04/02/04 13:25 SMT Quest rebooted a router and now the link to Mt Graham appears to be back to full speed. Fixed
04/02/04 13:15 12m ssh version 3.8pl is now installed at Kitt Peak and appears to be functioning correctly. News
04/02/04 09:00 12m We are in the process of upgrading our ssh installation so during your next login to any of our computers, you may get a warning about host-key being different. This is normal and you are safe to ignore and update your host-keys. News
04/01/04   SMT The 345 SIS receiver has been pulled and move to the downtown offices for repairs.  
04/01/04 12:00 SMT It has been reported that the network bandwidth to and from Mt. Graham is flakey at best. The start of this problem is coincident with the installation of a new router at Center for Computing and Information Technology, CCIT, on March 22. Ongoing
04/01/04 11:00 12m Operator's outgoing mail is now working. Fixed
04/01/04 9:30 12m Due to failure of modelo, email to and from the mountain operator has been interrupted. Fixed

 Copyright Arizona Radio Observatory.
For problems or questions regarding this web contact [tfolkers{at}email{dot}arizona{dot}edu].
Last updated: 05/18/11.